A bag contains 10 red balls, 30 white balls, 20 black balls, and 15 yellow balls. If 2 balls are drawn, replacement being made after each drawing, find the probability that only one is red.
1
Expert's answer
2020-03-03T10:04:37-0500

Solution. When drawing two balls, the red ball can be selected from both the first and second attempts. Therefore the probability is equal to


"p=\frac{65}{75}\times\frac{10}{75}+\frac{10}{75}\times\frac{65}{75}=\frac{1300}{5625}\approx 0.2311"

Answer. 0.2311